OpenStudy (anonymous):

WILL MEDAL!! Find the equation of the axis of symmetry and the coordinates of the vertex of the graph of each function. 7. y=2x^2+3 9. f(x)=2x^2+4x-1 11. f(x)=3x^2-9x+2 13. f(x)=-5x^2+3x+2 15. f(x)=6x^2+6x-5

OpenStudy (anonymous):

@Phebe

OpenStudy (phebe):

7. -2 x^2+y-3 = 0

OpenStudy (phebe):

9. f(x) = 2 x (x+2)-1

OpenStudy (phebe):

11. 3 (x-3) x+2 = f(x)

OpenStudy (phebe):

13. f(x) = (3-5 x) x+2

OpenStudy (phebe):

15. f(x) = 6 x (x+1)-5

OpenStudy (phebe):

There yah go!! :D

OpenStudy (anonymous):

Thank you so much @Phebe !!!